In an arithmetic series the sum of the first 9 terms is 162 and the sum of the first 12 terms is 288. Determine the series.
Answer by Theo(3464)

sum of the terms of an arithmetic series is given by the formula:

s[n] = n * (a[1] + a[n])/2

s[n] = sum of the n terms of the sequence.
a[1] = first term in the sequence.
a[n] = nth term in the sequence.
n = number of terms in the sequence.

the formuula for the nth term in the sequence is:

a[n] = a[1] + (n-1) * d

d is the common difference.

you will use both these formulas to find the answer.

you are given:

sum of the first 9 terms in a sequence is 162
sum of the first 12 terms in the same sequence is 288.

the formnula for the sum of the first n terms in a sequence is:

s[n] = n * (a[1] + a[n])/2

substitute for the first 9 terms to get:

162 = 9 * (a[1] + a[9])/2 (first equation)

substitute for the first 12 terms to get:

288 = 12 * (a[1] + a[12])/2 (second equation)

multiply both sides of the first equation by 2 and divide both sides of the first equation by 9 to get:

a[1] + a[9] = 36 (third equation)

multiply both sides of the second equation by 2 and divide both sides of the second equation by 12 to get:

a[1] + a[12] = 48 (fourth equation)

subtract third equation from fourth equation to get:

a[12] - a[9] = 12

since 12 - 9 = 3, there are 3 terms between a[9] and a[12].

12/3 = 4 which means the common difference between each term is 4.

the nth term in each sequence is given by the formula:

a[n] = a[1] + (n-1) * d

we now know that d = 4

we solve for a[9] to get:

a[9] = a[1] + 8*4 = a[1] + 32

we take our first equation and plug in our new found values.

first equation is:

162 = 9 * (a[1] + a[9])/2 (first equation)

substitute a[1] + 32 for a[9] to get:

162 = 9 * (a[1] + a[1] + 32)/2

combine like terms to get:

162 = 9 * (2* a[1] + 32)/2

multiply both sides of this equation by 2 and divide both sides of this equation by 9 to get:

36 = 2 * a[1] + 32

subtract 32 from both sides of this equation to get:

36-32 = 2 * a[1] which becomes:

4 = 2 * a[1].

divide both sides of this equation by 2 to get:

a[1] = 2.

we now know a[1] and we know d so we can construct the sequence.

formula for nth term in the sequence is:

a[n] = a[1] + (n-1) * d

since d = 4,...

the 9th term in the sequence is:

a[9] = 2 + 8*4 = 2 + 32 = 34

the 12th term in the sequence is:

a[12] = 2 + 11*4 = 2 + 44 = 46

the sequence is:
         
term:         1  2  3   4   5   6   7   8   9   10  11  12

sequence:     2  6  10  14  18  22  26  30  34  38  42  46

2 is the 1st term.
34 is the 9th term.
46 is the 12th term.

common difference is 4.